We have replaced faulty equipment. We've now resolved the incident. Thanks for your patience.


The switching hardware we replaced yesterday was apparently not the root cause. The defect is more significant. We are working with our providers to get an emergency delivery of replacement equipment to get this back up and operational. We may have a temporary spare to utilize to make this happen quicker.

Avatar for

The Luth complex is again having issues in certain areas this morning. We're working to resolve this with our partners.

Avatar for
Began at:

Affected components
  • The Internet
  • Wireless Network
  • Wired Network